OpenStudy (anonymous):

if three times a number added to 2 is divided by the number plus 5, the result is five halves, find the number

OpenStudy (anonymous):

if I understand correctly, the equation should look like this: \[\frac{ 3x + 2 }{ x + 5 } = \frac{ 5 }{ 2 }\] now we multiply denominator of the left fraction with numerator of the right fraction and vice verse \[(3x+2)*2=(x+5)*5\]\[6x + 4 = 5x + 25\] from here on, it shouldn't be a problem; however, if you can't find the solution, feel free to ask;)

OpenStudy (anonymous):

i cant find the solution =/

OpenStudy (anonymous):

one more hint, put the unknowns to left side and known values to the right side of the = sign